1) Suppose that we have a collection of brown and white eggs. Suppose 1/3 of the brown eggs are rotten, and 7/10 of the white eggs are rotten.

a) If an egg is chosen at random, with an equal chance of being white or brown, what is the probability that it is rotten?

b) Suppose that an egg is chosen at random, with an equal chance of being white or brown. If the egg is found to be rotten, what is the probability that it was a white egg?
Answer #1

P(R|B)=1/3

P(B)=1/2

P(W)=1/2

P(R|W)=7/10

a)

b)

P(W|R)

=0.6774